# OFFSET: Segmentation-based Focus Shift Revision for Composed Image Retrieval

### v1

@@ 4.2. Performance Comparison
We compare OFFSET with several CIR methods, which can be classified into two categories according to the utilized backbone: traditional model-based baselines (e.g., TIRG ((Vo et al., 2019)), MGUR ((Chen et al., 2024))) and CLIP-based baselines (e.g. DQU-CIR ((Wen et al., 2024)), ENCODER ((Li et al., 2025))). Our analysis of the comparative data in Table 4.1.3, Table 4.1.3, and Table 4.1.3 yields the following significant observations: 1) OFFSET consistently outperforms all baseline models on FashionIQ, Shoes, and CIRR datasets. Specifically, OFFSET achieves $1.00$% relative improvements over the best baseline for R@$10$ on FashionIQ-Avg, $0.89$% for Rsubset@$1$ on CIRR, and $0.79$% for average on Shoes, respectively, demonstrating its effectiveness and generalization ability in both fashion-specific and open-domain CIR tasks. 2) The VLP-based models (bottom of tables) typically outperform those based on traditional feature extraction models (top of tables), which confirms the effectiveness of VLP-based models applied to the CIR task and provides a solid foundation for visual-textual semantic alignment. 3) DQU-CIR outperforms all other baselines on all metrics in the fashion-domain dataset, but its performance on the open-domain dataset CIRR remains inferior. This may be due to the complexity of the open-domain dataset, which hinders the model’s OCR capability to accurately recognize the semantics of multimodal queries’ keywords, thus limiting its performance. In contrast, OFFSET not only outperforms the previous models in fashion-domain datasets but also achieves state-of-the-art results on the open-domain dataset CIRR, which indicates that OFFSET has stable CIR performance and is not limited to domain-specific data.
@@ 4.3. Ablation Studies
From the results of G1 in Table 4.1.3, we can obtain the following observations. 1) D#(1) w/o FM performs the worst among the variants in this group. This demonstrates the necessity of simultaneously enabling the visual and textual features to focus on the portions that are closely related to the multimodal queries. 2) both D#(2) and D#(3) are inferior to OFFSET, which indicates that independently performing focus mapping for visual and textual features can also effectively focus their portions close to the semantics closely related to the multimodal queries. 3) The degradation in performance of D#(4) reveals the importance of Multi-Grained Focus Projection in precisely aligning the local and global focus.
@@
From the experimental results of G2 in Table 4.1.3, the following findings can be obtained. 1) both D#(5) and D#(6) demonstrate a performance decline compared to the complete OFFSET. This indicates the necessity of performing Visual Focus Mapping and Multi-Grained Focus Projection on the target image, which makes the target semantics of focused portions more close to the semantics of multimodal queries. 2) D#(7) exhibits a significant gap compared to OFFSET, demonstrating that the focus revision process is indeed effective for revising the focus of the multimodal query.
$\bullet$ G3:Ablation on Optimization Functions This group is devised to explore the effect of the optimization functions of OFFSET, whose derivatives are listed as follows. D#(8) w/o BBC: To check the effect of the batch-based classification loss (BBC, Eqn.((((12))))), we remove $\mathcal{L}_{rank}$ in Eqn.((((15)))). D#(9) w/o FR: To validate the impact of focus regularization (FR, Eqn.((((14))))), we ablate $\mathcal{L}_{fr}$ in Eqn.((((15)))).
We can observe from the results of G3 in Table 4.1.3 that: 1) D#(8) performs worse than OFFSET, which proves the effectiveness of the batch-based classification loss in guiding the model to learn better multimodal focus features. 2) The performance drop of D#(9) indicates that focus regularization plays a vital role in maintaining focus consistency between the multimodal query and target.
@@ 4.4.2. Efficiency Analysis
In Table 5, we present a comparison of the inference efficiency between the proposed OFFSET and the representative CIR model DQU-CIR (Wen et al., 2024). Specifically, we list the inference time per sample, the corresponding retrieval performance on FashionIQ and CIRR, and additional auxiliary models used by the two models (i.e., the caption model (BLIP-2 (Li et al., 2023) for both DQU-CIR and OFFSET), the segmentation model (CLIPSeg (Lüddecke and Ecker, 2022) for OFFSET), and the LLM (Gemini-pro-v1 for DQU-CIR)). All experiments are conducted on a single A40 GPU. From the results in the table, we observe that the inference time of our proposed OFFSET decreases by $40.27$% compared to DQU-CIR, while the retrieval performance on all datasets outperforms that of DQU-CIR (as illustrated in Table 4.1.3, Table 4.1.3, and Table 4.1.3), especially with an improvement of $9.98$% on CIRR-Avg. This indicates that OFFSET achieves optimal retrieval performance without incurring excessive additional overhead, which aligns with the retrieval intent of the CIR task.
Table 5. Comparison of inference efficiency. The better results are in bold. Cap. represents the utilized caption model and Seg. denotes the utilized segmentation model. F-R10-Avg. is the average of R@$10$ on FashionIQ and C-Avg. represents the ((R@$5$+Rsubset@$1$)/$2$) on CIRR.
@@ 5. Conclusion
This work found that the CIR community has two phenomena, which is seriously neglected: 1) the inhomogeneity in visual data leads to the degradation of query features. 2) the text-priority in multimodal queries leads to visual focus bias. In light of these findings, we proposed OFFSET, which encompasses three advantages. Firstly, to address the inhomogeneity, we developed a focus mapping-based feature extractor, which identifies the dominant region and guides the visual and textual feature extraction, thereby reducing noise interference. Secondly, for the phenomenon of text-priority, we proposed textually guided focus revision, which enables the adaptive focus revision on the reference image according to the modification semantics, thus enhancing the perception of the modification focus on the composed features. Ultimately, extensive experiments on four benchmark datasets substantiated the efficacy of our proposed OFFSET. In the future, we intend to extend our approach to address other downstream tasks, such as information detection and prediction (Wu et al., 2025; Qiu et al., 2024; Liu et al., 2025; Qiu et al., 2025).

### v2

@@ 4.2. Performance Comparison
We compare OFFSET with several CIR methods, which can be classified into two categories according to the utilized backbone: traditional model-based baselines (e.g., TIRG ((Vo et al., 2019b)), MGUR ((Chen et al., 2024d))) and CLIP-based baselines (e.g. DQU-CIR ((Wen et al., 2024)), ENCODER ((Li et al., 2025b))). Our analysis of the comparative data in Table 1, Table 2, and Table 3 yields the following significant observations: 1) OFFSET consistently outperforms all baseline models on FashionIQ, Shoes, and CIRR datasets. Specifically, OFFSET achieves 1.001.00% relative improvements over the best baseline for R@$10$ on FashionIQ-Avg, 0.890.89% for Rsubset@$1$ on CIRR, and 0.790.79% for average on Shoes, respectively, demonstrating its effectiveness and generalization ability in both fashion-specific and open-domain CIR tasks. 2) The VLP-based models (bottom of tables) typically outperform those based on traditional feature extraction models (top of tables), which confirms the effectiveness of VLP-based models applied to the CIR task and provides a solid foundation for visual-textual semantic alignment. 3) DQU-CIR outperforms all other baselines on all metrics in the fashion-domain dataset, but its performance on the open-domain dataset CIRR remains inferior. This may be due to the complexity of the open-domain dataset, which hinders the model’s OCR capability to accurately recognize the semantics of multimodal queries’ keywords, thus limiting its performance. In contrast, OFFSET not only outperforms the previous models in fashion-domain datasets but also achieves state-of-the-art results on the open-domain dataset CIRR, which indicates that OFFSET has stable CIR performance and is not limited to domain-specific data.
@@ 4.3. Ablation Studies
From the results of G1 in Table 4, we can obtain the following observations. 1) D#(1) w/o FM performs the worst among the variants in this group. This demonstrates the necessity of simultaneously enabling the visual and textual features to focus on the portions that are closely related to the multimodal queries. 2) both D#(2) and D#(3) are inferior to OFFSET, which indicates that independently performing focus mapping for visual and textual features can also effectively focus their portions close to the semantics closely related to the multimodal queries. 3) The degradation in performance of D#(4) reveals the importance of Multi-Grained Focus Projection in precisely aligning the local and global focus.
@@
From the experimental results of G2 in Table 4, the following findings can be obtained. 1) both D#(5) and D#(6) demonstrate a performance decline compared to the complete OFFSET. This indicates the necessity of performing Visual Focus Mapping and Multi-Grained Focus Projection on the target image, which makes the target semantics of focused portions more close to the semantics of multimodal queries. 2) D#(7) exhibits a significant gap compared to OFFSET, demonstrating that the focus revision process is indeed effective for revising the focus of the multimodal query.
$\bullet$ G3:Ablation on Optimization Functions This group is devised to explore the effect of the optimization functions of OFFSET, whose derivatives are listed as follows. D#(8) w/o BBC: To check the effect of the batch-based classification loss (BBC, Eqn.((12))), we remove $\mathcal{L}_{rank}$ in Eqn.((15)). D#(9) w/o FR: To validate the impact of focus regularization (FR, Eqn.((14))), we ablate $\mathcal{L}_{fr}$ in Eqn.((15)).
We can observe from the results of G3 in Table 4 that: 1) D#(8) performs worse than OFFSET, which proves the effectiveness of the batch-based classification loss in guiding the model to learn better multimodal focus features. 2) The performance drop of D#(9) indicates that focus regularization plays a vital role in maintaining focus consistency between the multimodal query and target.
@@ 4.4.2. Efficiency Analysis
In Table 5, we present a comparison of the inference efficiency between the proposed OFFSET and the representative CIR model DQU-CIR (Wen et al., 2024). Specifically, we list the inference time per sample, the corresponding retrieval performance on FashionIQ and CIRR, and additional auxiliary models used by the two models (i.e., the caption model (BLIP-2 (Li et al., 2023) for both DQU-CIR and OFFSET), the segmentation model (CLIPSeg (Lüddecke and Ecker, 2022) for OFFSET), and the LLM (Gemini-pro-v1 for DQU-CIR)). All experiments are conducted on a single A40 GPU. From the results in the table, we observe that the inference time of our proposed OFFSET decreases by 40.2740.27% compared to DQU-CIR, while the retrieval performance on all datasets outperforms that of DQU-CIR (as illustrated in Table 1, Table 2, and Table 3), especially with an improvement of 9.989.98% on CIRR-Avg. This indicates that OFFSET achieves optimal retrieval performance without incurring excessive additional overhead, which aligns with the retrieval intent of the CIR task.
Table 5. Comparison of inference efficiency. The better results are in bold. Cap. represents the utilized caption model and Seg. denotes the utilized segmentation model. F-R10-Avg. is the average of R@$10$ on FashionIQ and C-Avg. represents the ((R@55+Rsubset@11)/22) on CIRR.
@@ 5. Conclusion
This work found that the CIR community has two phenomena, which is seriously neglected: 1) the inhomogeneity in visual data leads to the degradation of query features. 2) the text-priority in multimodal queries leads to visual focus bias. In light of these findings, we proposed OFFSET, which encompasses three advantages. Firstly, to address the inhomogeneity, we developed a focus mapping-based feature extractor, which identifies the dominant region and guides the visual and textual feature extraction, thereby reducing noise interference. Secondly, for the phenomenon of text-priority, we proposed textually guided focus revision, which enables the adaptive focus revision on the reference image according to the modification semantics, thus enhancing the perception of the modification focus on the composed features. Ultimately, extensive experiments on four benchmark datasets substantiated the efficacy of our proposed OFFSET. In the future, we intend to extend our approach to address other downstream tasks, such as information detection and prediction (Wu et al., 2025c; Qiu et al., 2024; Liu et al., 2025c; Qiu et al., 2025c).